Pecan Theatre Inc. owns and operates movie theaters throughout Florida and Georgia. Pecan Theatre has declared the following annual dividends over a six-year period: Year 1, $20,000; Year 2, $60,000; Year 3, $90,000; Year 4, $110,000; Year 5, $140,000; and Year 6, $180,000. During the entire period ending December 31 of each year, the outstanding stock of the company was composed of 25,000 shares of cumulative, 2% preferred stock, $100 par, and 100,000 shares of common stock, $15 par.

Required:

1. Determine the total dividends and the per-share dividends declared on each class of stock for each of the six years. There were no dividends in arrears at the beginning of Year 1. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0".

Preferred Dividends Common Dividends

Year
Total
Dividends

Total

Per Share

Total

Per Share
Year 1 $   20,000 $ $ $ $
Year 2 60,000
Year 3 90,000
Year 4 110,000
Year 5 140,000
Year 6 180,000
$ $

2. Determine the average annual dividend per share for each class of stock for the six-year period. If required, round your answers to two decimal places.

Average annual dividend for preferred $ per share
Average annual dividend for common $ per share

3. Assuming a market price per share of $159 for the preferred stock and $20 for the common stock, determine the average annual percentage return on initial shareholders' investment, based on the average annual dividend per share for preferred stock and for common stock.

Round your answers to two decimal places.

Preferred stock %
Common stock %
Answer #1

1. Determine the total dividends and the per-share dividends declared on each class of stock for each of the six years. There were no dividends in arrears at the beginning of Year 1. Summarize the data in tabular form. If required, round your answers to two decimal places. If the amount is zero, please enter "0".

Preferred dividend = 25000*100*2% = 50000

Preferred Dividends Common Dividends

Year
Total
Dividends

Total

Per Share

Total

Per Share
Year 1 $   20,000 $20000 $0.80 $0 $0
Year 2 60,000 60000 2.40 0 0
Year 3 90,000 70000 2.80 20000 0.20
Year 4 110,000 50000 2.00 60000 0.60
Year 5 140,000 50000 2.00 90000 0.90
Year 6 180,000 50000 2.00 130000 1.30
600000 300000 $12 300000 $3.00

2. Determine the average annual dividend per share for each class of stock for the six-year period. If required, round your answers to two decimal places.

Average annual dividend for preferred $12/6 = 2 per share
Average annual dividend for common $3/6 = 0.50 per share

3. Assuming a market price per share of $159 for the preferred stock and $20 for the common stock, determine the average annual percentage return on initial shareholders' investment, based on the average annual dividend per share for preferred stock and for common stock.

Round your answers to two decimal places.

Preferred stock 2/159 = 1.26%
Common stock 0.50/20 = 2.5%